# Rethinking Aleatoric and Epistemic Uncertainty

### v2

@@ 1 Introduction
Decision problem $\,\ell:\mathcal{A}\times\mathcal{Z}\to\mathbb{R}$ Training data $\,y_{1:n}\sim p_{\mathrm{train}}(y_{1:n}|\pi)$ Predictive model $\,p_{n}(z)=p(z;y_{1:n})$ Bayes-optimal action $\,a_{n}^{*}=\arg\min\nolimits_{a\in\mathcal{A}}\mathbb{E}_{p_{n}(z)}\!\left[\ell(a,z)\right]$ Predictive uncertainty $\,h\!\left[p_{n}(z)\right]=\mathbb{E}_{p_{n}(z)}[\ell(a_{n}^{*},z)]$ $\,\mathrm{EUR}_{z}^{\mathrm{true}}(\pi,m)$ $=h\!\left[p_{n}(z)\right]-\mathbb{E}_{p_{\mathrm{train}}(y^{\text{\scriptsize+}}_{1:m}|\pi)}\!\left[h\!\left[p_{n+m}(z)\right]\right]$ $\approx h\!\left[p_{n}(z)\right]-\mathbb{E}_{p_{n}(y^{\text{\scriptsize+}}_{1:m}|\pi)}\!\left[h\!\left[q_{n+m}(z)\right]\right]$ Expected uncertainty reduction Uncertainty Proper scoring rule $\,s(p_{n},z)=\ell(a^{*}_{n},z)$ Discrepancy function $\,d(p_{n},p_{\mathrm{eval}})=\mathbb{E}_{p_{\mathrm{eval}}(z)}\!\left[s(p_{n},z)\right]-h\!\left[p_{\mathrm{eval}}(z)\right]$ Evaluation LearningReasoning
Figure 1: Our decision-theoretic view allows us to disentangle machine-learning concepts that have been conflated under the aleatoric-epistemic view. We consider taking an action, $a\in\mathcal{A}$, in light of imperfect knowledge of $z\in\mathcal{Z}$, with an action’s consequences measured by a loss function, $\ell(a,z)$. Since $z$ is unknown, we use any available training data, $y_{1:n}$, to build a predictive model, $p_{n}(z)$, with which we can reason over possible values of $z$ and thus choose an action. Additionally we can use the model to quantify uncertainty and its reducibility with respect to new data, and we can evaluate the model using a ground-truth value of $z$ or a reference distribution, $p_{\mathrm{eval}}(z)$.
When making decisions under uncertainty, it can be useful to reason about where that uncertainty comes from (Osband et al,, 2023; Wen et al,, 2022). Researchers commonly refer to the ideas of aleatoric (literal meaning: “relating to chance”) and epistemic (“relating to knowledge”) uncertainty, which have a long history in the study of probability (Hacking,, 1975). Aleatoric uncertainty is typically associated with statistical dispersion in data (sometimes thought of as noise), while epistemic is associated with the internal information state of a model (Hüllermeier & Waegeman,, 2021).
@@
To establish a clearer perspective, we draw on powerful yet underappreciated ideas from decision theory (Dawid,, 1998; DeGroot,, 1962; Neiswanger et al,, 2022). Our starting point is a final decision of interest with an associated loss function. Given this, uncertainty in predictive beliefs can be formalised as the subjective expected loss of acting Bayes-optimally under those beliefs; this generalises quantities like variance and entropy. From there we show how reasoning about new data gives rise to a notion of expected uncertainty reduction, which we can use to identify a decomposition of uncertainty into irreducible and reducible components. Then we clarify the connection between uncertainty, predictive performance and data dispersion, linking to classic decompositions from statistics and information theory. Overall this provides a coherent synthesis of key quantities that researchers are interested in (Figure 1).
@@
Our work thus serves to inform future work in two key ways. On the one hand it sheds light on the contradictions of the aleatoric-epistemic view and presents a coherent alternative perspective that allows clearer thinking about uncertainty in machine learning. On the other hand it provides more direct practical insights. It clarifies that what might have seemed like arbitrary choices for a decision-maker can instead be made by following well-defined logic: given some basic components and principles, it becomes clear how we should measure predictive uncertainty, how to assess models using reference systems and how to identify good future training data. It also highlights approximations that often have to be made in practice, revealing scope for suboptimal performance and therefore informing future methods research.
@@ 2 Background
The broad idea of the aleatoric-epistemic decomposition is to distinguish between different sources of uncertainty. If a model’s prediction is uncertain, we might want to know whether that prediction is fundamentally uncertain for the given model class or instead due to a lack of data. This breakdown has clear utility in the context of seeking new data that will reduce predictive uncertainty (Bickford Smith et al,, 2023, 2024; MacKay, 1992a, ; MacKay, 1992b, ). But it is also relevant elsewhere: in model selection, for example, we might want to quantify a model’s scope for improvement by forecasting how its predictions will change given more data (Barbieri & Berger,, 2004; Fong & Holmes,, 2020; Geisser & Eddy,, 1979; Kadane & Lazar,, 2004; Laud & Ibrahim,, 1995).
@@ 3 Key concepts
Our aim in this work is to formalise and unify quantities that have been associated with the ideas of aleatoric and epistemic uncertainty in past work. In particular we look to identify a rigorous notion of predictive uncertainty and the extent to which it reduces as more data is observed, and also measures of predictive performance and statistical dispersion in data. We start by highlighting some foundational concepts that will be used throughout our discussion.
@@ 3.3 Subjective expected loss enables decision-making
Since $\ell$ is a function of the unknown $z$, it cannot be used directly as an objective for selecting an action, $a$. A principled solution that we focus on here is to form subjective beliefs over possible values of $z$ (conventionally this belief state would be a Bayesian prior or posterior), average over these to form an expected loss, then choose an action that minimises this subjective expected loss (Ramsey,, 1926; Savage,, 1951). Alternative approaches to decision-making include minimax (von Neumann,, 1928; Wald,, 1949), which involves acting so as to minimise the worst-case loss.
@@ 3.5 Bayes optimality is a subjective notion
“captures noise inherent in the observations” $\,\mathrm{H}\!\left[p_{\mathrm{train}}(y_{i}|\pi(y_{<i}),y_{<i})\right]$ or $\mathrm{H}\!\left[p_{\mathrm{eval}}(z)\right]$“cannot be reduced even if more data wereto be collected” $\,\mathrm{H}\!\left[p_{\infty}(z)\right]$Expected conditional predictive entropy $\,\mathbb{E}_{p_{n}(\theta)}\!\left[\mathrm{H}\!\left[p_{n}(z|\theta)\right]\right]$123Aleatoric uncertainty “uncertainty in the model parameters” $\,\mathrm{H}\!\left[p_{n}(\theta)\right]$“can be explained away given enough data” $\,\mathrm{H}\!\left[p_{n}(z)\right]-\mathrm{H}\!\left[p_{\infty}(z)\right]$Expected information gain in the parameters $\,\mathrm{H}\!\left[p_{n}(z)\right]-\mathbb{E}_{p_{n}(\theta)}\!\left[\mathrm{H}\!\left[p_{n}(z|\theta)\right]\right]$123Epistemic uncertainty
Figure 2: A popular view on aleatoric and epistemic uncertainty in machine learning attaches multiple mathematical quantities to each of the two concepts, rendering it incoherent and thus a likely source of conflations in the literature. Some of these quantities can coincide in particular cases but in the general case they are distinct. The quotations here are from Kendall & Gal, (2017) and have been expressed mathematically as information-theoretic quantities. The interpretation of Equation 1 is due to Gal, (2016) and Gal et al, (2017).
@@ 3.6 Predictions often do not match data generation
3.7 Reference systems allow grounded evaluation
Because we expect $p_{n}(z)$ to be imperfect, we typically want to evaluate it against a reference system, $p_{\mathrm{eval}}(z)$, which could for example be a computer program, a human expert or a physical sensor. It is common in machine learning to evaluate models using an estimator of the frequentist risk (Berger,, 1985), such as the mean squared error on finite data sampled from $p_{\mathrm{eval}}(z)$. Notably $p_{\mathrm{eval}}(z)$ could itself be imperfect (Fluri et al,, 2023), so care is needed in designing and interpreting evaluations using reference systems.
@@ 4 Assessing a popular view
[…]
While that work successfully captured some of the intuitions from Section 2, we highlight that it also overloaded the ideas of aleatoric and epistemic uncertainty with multiple meanings, introducing a number of spurious associations (Figure 2). The competing definitions of aleatoric uncertainty conflate $\mathrm{H}\!\left[p_{\infty}(z)\right]$, measuring the uncertainty in $p_{n}(z)$ as $n\to\infty$ (this depends on the data-generating process, as we will discuss in Section 5.2), with three separate quantities:
@@
    $\mathrm{H}\!\left[p_{\mathrm{train}}(y_{i}|\pi(y_{<i}),y_{<i})\right]$, the entropy in training-data generation. Issue: $p_{\infty}(z)$ is a subjective belief state that need not match $p_{\mathrm{train}}(y_{i}|\pi(y_{<i}),y_{<i})$ (Section 3.6).
@@
    $\mathrm{H}\!\left[p_{\mathrm{eval}}(z)\right]$, the entropy of the reference system used in evaluation. Issue: $p_{\infty}(z)$ is a subjective belief state that need not match $p_{\mathrm{eval}}(z)$ (Sections 3.6 and 3.7).
@@ 5.1 Predictive uncertainty can be derived from the final decision of interest and the associated loss
    $h\!\left[p_{n}(z)\right]$ $=\mathbb{E}_{p_{n}(z)}\!\left[(\mathbb{E}_{p_{n}(z)}\!\left[z\right]-z)^{2}\right]=\mathbb{V}_{p_{n}(z)}\!\left[z\right].\qed$
@@
    $h\!\left[p_{n}(z)\right]$ $=-\mathbb{E}_{p_{n}(z)}\!\left[\log p_{n}(z)\right]=\mathrm{H}\!\left[p_{n}(z)\right].\qed$
@@ 5.2 Decomposing predictive uncertainty requires accounting for the data-generating process
Characterising the reducibility of uncertainty initially seems as simple as considering new data, $y^{\text{\scriptsize+}}_{1:m}=y_{(n+1):(n+m)}$, and measuring the corresponding uncertainty reduction,
$\mathrm{UR}_{z}(y^{\text{\scriptsize+}}_{1:m})=h\!\left[p_{n}(z)\right]-h\!\left[p_{n+m}(z)\right],$
which we note could be negative. But this uncertainty reduction depends on exactly what the new data is, and that in turn depends on the process by which the data is generated. We therefore need to explicitly account for the data-generating process to produce a well-defined notion of reducibility. Any stochasticity in model updating also has to be taken into account, but this can be achieved by absorbing the updating stochasticity into the definition of $y^{\text{\scriptsize+}}_{1:m}$ (Section 3.4).
Revisiting $p_{\mathrm{train}}(y^{\text{\scriptsize+}}_{1:m}|\pi)$ from Section 3.4, we define the distribution over $y^{\text{\scriptsize+}}_{i}$ to depend both on decisions made by the data-acquisition policy, $\pi$, and on the previous data:
$p_{\mathrm{train}}(y^{\text{\scriptsize+}}_{1:m}|\pi)=\prod_{i=1}^{m}p_{\mathrm{train}}(y^{\text{\scriptsize+}}_{i}|\pi(y_{<i}),y_{<i}).$
@@
$\mathrm{EUR}_{z}^{\mathrm{true}}(\pi,m)=\mathbb{E}_{p_{\mathrm{train}}(y^{\text{\scriptsize+}}_{1:m}|\pi)}\!\left[\mathrm{UR}_{z}(y^{\text{\scriptsize+}}_{1:m})\right].$ (3)
@@
$\underbrace{\mathrm{EUR}_{z}^{\mathrm{true}}(\pi,\infty)}_{\mathrm{reducible}}=\underbrace{h\!\left[p_{n}(z)\right]}_{\mathrm{total}}-\underbrace{\mathbb{E}_{p_{\mathrm{train}}(y^{\text{\scriptsize+}}_{1:\infty}|\pi)}\!\left[h\!\left[p_{\infty}(z)\right]\right]}_{\mathrm{irreducible}}.$
@@
It is worth noting that in some restricted cases the dependency on the data-acquisition policy, $\pi$, of the infinite-data terms in this uncertainty decomposition can disappear. For example, if we are in a supervised-learning setting where the policy’s decisions are based on inputs to acquire labels for, and if we are using a well-specified Bayesian model and exact Bayesian updating, $p_{\infty}(z)$ should be independent of $\pi$ as long as $\pi$ produces dense samples across the input space (Kleijn & van der Vaart,, 2012). However, the requirements for this are very strict, with any model misspecification or error in belief updating reintroducing the dependency.

### v3

@@ 1 Introduction
Decision problem $\ell:\mathcal{A}\times\mathcal{Z}\to\mathbb{R}$ Training data $y_{1:n}\sim p_{\mathrm{train}}(y_{1:n}|\pi)$ Predictive model $p_{n}(z)=p(z;y_{1:n})$ Bayes-optimal action $a_{n}^{*}=\arg\min\nolimits_{a\in\mathcal{A}}\mathbb{E}_{p_{n}(z)}\!\left[\ell(a,z)\right]$ Predictive uncertainty $h\!\left[p_{n}(z)\right]=\mathbb{E}_{p_{n}(z)}[\ell(a_{n}^{*},z)]$ Expected uncertainty reduction $\,\mathrm{EUR}_{z}^{\mathrm{true}}(\pi,m)$ $=h\!\left[p_{n}(z)\right]-\mathbb{E}_{p_{\mathrm{train}}(y^{\scriptscriptstyle+}_{1:m}|\pi)}\!\left[h\!\left[p_{n+m}(z)\right]\right]$ $\approx h\!\left[p_{n}(z)\right]-\mathbb{E}_{p_{n}(y^{\scriptscriptstyle+}_{1:m}|\pi^{\prime})}\!\left[h\!\left[q_{n+m}(z)\right]\right]$ Uncertainty quantification Proper scoring rule $s(p_{n},z)=\ell(a^{*}_{n},z)$ Discrepancy function $d(p_{n},p_{\mathrm{eval}})=\mathbb{E}_{p_{\mathrm{eval}}(z)}\!\left[s(p_{n},z)\right]-h\!\left[p_{\mathrm{eval}}(z)\right]$ Model evaluation LearningReasoning
Figure 1: Our decision-theoretic view coherently relates machine-learning concepts that have been conflated under the aleatoric-epistemic view. We consider taking an action, $a\in\mathcal{A}$, in light of imperfect knowledge of $z\in\mathcal{Z}$, with an action’s consequences measured by a loss function, $\ell(a,z)$. Since $z$ is unknown, we use any available training data, $y_{1:n}$, to build a predictive model, $p_{n}(z)$, with which we can reason over possible values of $z$ and thus choose an action. Additionally we can use the model to quantify uncertainty and its reducibility with respect to new data, and we can evaluate the model using a ground-truth realisation of $z$ or a reference distribution, $p_{\mathrm{eval}}(z)$.
When making decisions under uncertainty, it can be useful to reason about where that uncertainty comes from (Osband et al,, 2023; Wen et al,, 2022). Researchers often aim to do this by referring to aleatoric (literal meaning: “relating to chance”) and epistemic (“relating to knowledge”) uncertainty, ideas with a long history in the study of probability (Hacking,, 1975). Aleatoric uncertainty is typically associated with statistical dispersion in data (sometimes thought of as noise), while epistemic is associated with a model’s internal information state (Hüllermeier & Waegeman,, 2021).
@@
To establish a clearer perspective, we draw on powerful yet underappreciated ideas from decision theory (Dawid,, 1998; DeGroot,, 1962; Neiswanger et al,, 2022). Our starting point is a final decision of interest with an associated loss function. Given this, uncertainty in predictive beliefs can be formalised as the subjective expected loss of acting Bayes-optimally under those beliefs; this generalises quantities like variance and Shannon entropy. From there we show how reasoning about new data gives rise to a notion of expected uncertainty reduction, which we can use to identify a decomposition of uncertainty into irreducible and reducible components. Then we clarify the connection between uncertainty, predictive performance and data dispersion, linking to classic decompositions from statistics and information theory. Overall this provides a coherent synthesis of key quantities that researchers are interested in (Figure 1).
@@
Our work thus serves to inform future work in two key ways. On the one hand it sheds light on the contradictions of the aleatoric-epistemic view and presents a coherent alternative perspective that allows clearer thinking about uncertainty in machine learning. On the other hand it provides more direct practical insights. It clarifies that what might have seemed like arbitrary choices for a decision-maker can instead be made by following well-defined logic: given some basic components and principles, it becomes clear how we should measure predictive uncertainty, predictive performance and data dispersion, and how we can identify good future training data. It also highlights approximations that often have to be made in practice, revealing scope for suboptimal performance and therefore informing future methods research.
@@ 2 Background
The broad motivation of the aleatoric-epistemic view is to distinguish between different sources of uncertainty. If a model’s prediction is uncertain, we might want to know whether that prediction is fundamentally uncertain for the given model class or instead due to a lack of data. This breakdown has clear utility in the context of seeking new data that will reduce predictive uncertainty (Bickford Smith et al,, 2023, 2024; MacKay, 1992a, ; MacKay, 1992b, ). But it is also relevant elsewhere: in model selection, for example, we might want to quantify a model’s scope for improvement by forecasting how its predictions will change given more data (Barbieri & Berger,, 2004; Fong & Holmes,, 2020; Geisser & Eddy,, 1979; Kadane & Lazar,, 2004; Laud & Ibrahim,, 1995).
@@ 3 Key concepts
Our aim in this work is to formalise and link together quantities that have been associated with the ideas of aleatoric and epistemic uncertainty in past work. In particular we look to identify a rigorous notion of predictive uncertainty and the extent to which it reduces as more data is observed, and also measures of predictive performance and statistical dispersion in data. We start by highlighting some foundational concepts that will be used throughout our discussion.
@@ 3.3 Subjective expected loss enables decision-making
Since $\ell$ is a function of the unknown $z$, it cannot be used directly as an objective for selecting an action, $a$. A principled solution that we focus on here is to form subjective beliefs over possible values of $z$ (conventionally this belief state would be a Bayesian prior or posterior), average over these to form an expected loss, then choose an action that minimises this subjective expected loss (Ramsey,, 1926; Savage,, 1951). Alternative decision-making approaches include minimax, which involves minimising the worst-case frequentist risk (von Neumann,, 1928; Wald,, 1939, 1945).
@@ 3.5 Bayes optimality is a subjective notion
“captures noise inherent in the observations” $\mathrm{H}\!\left[p_{\mathrm{train}}(y_{1:n}|\pi)\right]$ or $\mathrm{H}\!\left[p_{\mathrm{eval}}(z)\right]$ “cannot be reduced even if more data were to be collected” $\mathrm{H}\!\left[p_{\infty}(z)\right]$ Expected parameter-conditional predictive entropy $\mathbb{E}_{p_{n}(\theta)}\!\left[\mathrm{H}\!\left[p_{n}(z|\theta)\right]\right]$ “uncertainty in the model parameters” $\mathrm{H}\!\left[p_{n}(\theta)\right]$ “can be explained away given enough data” $\mathrm{H}\!\left[p_{n}(z)\right]-\mathrm{H}\!\left[p_{\infty}(z)\right]$ Expected information gain in the model parameters $\mathrm{H}\!\left[p_{n}(z)\right]-\mathbb{E}_{p_{n}(\theta)}\!\left[\mathrm{H}\!\left[p_{n}(z|\theta)\right]\right]$ Aleatoric uncertainty Epistemic uncertainty $\neq\neq\neq\neq$
Figure 2: A popular interpretation of aleatoric and epistemic uncertainty in machine learning attaches multiple mathematical quantities to each of the two concepts, rendering it incoherent and thus a likely source of conflations in the literature. Here we show quotations from Kendall & Gal, (2017) expressed mathematically as information-theoretic quantities, along with an interpretation of Equation 1 due to Gal, (2016) and Gal et al, (2017). Some of the quantities shown can coincide in particular cases but in the general case they are distinct.
@@ 3.6 Predictions often do not match data generation
3.7 Model evaluation relies on external grounding
Because we expect $p_{n}(z)$ to be imperfect, we often want to assess it against a ground-truth realisation of $z$ or a reference distribution, $p_{\mathrm{eval}}(z)$. Commonly this distribution—which could for example represent a computer program, a human expert or a physical sensor—is used as source of evaluation data for computing an estimator of frequentist risk (Berger,, 1985), such as the mean squared error. Notably $p_{\mathrm{eval}}(z)$ could itself be imperfect (Fluri et al,, 2023), so designing and interpreting evaluations of this form requires care.
@@ 4 Assessing a popular view
where $\mathrm{H}$ denotes Shannon entropy and $\mathrm{EIG}_{\theta}$, also known as the BALD score (Houlsby et al,, 2011), is the expected information gain in $\theta$ from observing $z$. Gal, (2016) stated the $\text{total}=\text{aleatoric}+\text{epistemic}$” relationship and the correspondence between $p_{n}(z)$ and total uncertainty, while Gal et al, (2017) made the explicit link to Equation 1, informed by Houlsby et al, (2011). Kendall & Gal, (2017) discussed the aleatoric-epistemic view in the context of computer vision.
While that work successfully captured some of the intuitions from Section 2, we highlight that it also overloaded the ideas of aleatoric and epistemic uncertainty with multiple meanings, introducing a number of spurious associations (Figure 2). The competing definitions of aleatoric uncertainty conflate $\mathrm{H}\!\left[p_{\infty}(z)\right]$, the entropy of $p_{n}(z)$ as $n\to\infty$ (this depends on the data-generating process, as we will discuss in Section 5.2), with three separate quantities:
@@
    $\mathrm{H}\!\left[p_{\mathrm{train}}(y_{1:n}|\pi)\right]$, the entropy in training-data generation. Issue: $p_{\infty}(z)$ represents subjective predictive beliefs that need not match $p_{\mathrm{train}}(y_{1:n}|\pi)$ (Section 3.6).
@@
    $\mathrm{H}\!\left[p_{\mathrm{eval}}(z)\right]$, the entropy in evaluation-data generation. Issue: $p_{\infty}(z)$ represents subjective predictive beliefs that need not match $p_{\mathrm{eval}}(z)$ (Sections 3.6 and 3.7).
@@ 5.1 Predictive uncertainty can be derived from the final decision of interest and the associated loss
    $h\!\left[p_{n}(z)\right]$ $=\mathbb{E}_{p_{n}(z)}\!\left[(\mathbb{E}_{p_{n}(z)}\!\left[z\right]-z)^{2}\right]=\mathbb{V}_{p_{n}(z)}\!\left[z\right].$
@@
    $h\!\left[p_{n}(z)\right]$ $=-\mathbb{E}_{p_{n}(z)}\!\left[\log p_{n}(z)\right]=\mathrm{H}\!\left[p_{n}(z)\right].$
@@ 5.2 Decomposing predictive uncertainty requires accounting for the data-generating process
Characterising the reducibility of uncertainty initially seems as simple as considering new data, $y^{\scriptscriptstyle+}_{1:m}=y_{(n+1):(n+m)}$, and measuring the corresponding uncertainty reduction,
$\mathrm{UR}_{z}(y^{\scriptscriptstyle+}_{1:m})=h\!\left[p_{n}(z)\right]-h\!\left[p_{n+m}(z)\right],$
which we note could be negative. But this uncertainty reduction depends on exactly what the new data is, and that in turn depends on the process by which the data is generated. We therefore need to explicitly account for the data-generating process to produce a well-defined notion of reducibility. Any stochasticity in model updating also has to be taken into account, but this can be achieved by absorbing the updating stochasticity into the definition of $y^{\scriptscriptstyle+}_{1:m}$ (Section 3.4).
Revisiting the data-generating process from Section 3.4, we define the distribution over $y^{\scriptscriptstyle+}_{i}$ to depend on decisions made by the data-acquisition policy, $\pi$, and on the previous data:
$p_{\mathrm{train}}(y^{\scriptscriptstyle+}_{1:m}|\pi)=\prod_{i=1}^{m}p_{\mathrm{train}}(y^{\scriptscriptstyle+}_{i}|\pi(y_{<i}),y_{<i}).$
@@
[…]
@@
$\underbrace{\mathrm{EUR}_{z}^{\mathrm{true}}(\pi,\infty)}_{\mathrm{reducible}}=\underbrace{h\!\left[p_{n}(z)\right]}_{\mathrm{total}}-\underbrace{\mathbb{E}_{p_{\mathrm{train}}(y^{\scriptscriptstyle+}_{1:\infty}|\pi)}\!\left[h\!\left[p_{\infty}(z)\right]\right]}_{\mathrm{irreducible}}.$
@@
It is worth noting that in some restricted cases the dependency on the data-acquisition policy, $\pi$, of the infinite-data terms in this uncertainty decomposition can disappear. For example, if we are in a supervised-learning setting where the policy’s decisions concern which inputs to acquire labels for, and if we are using a well-specified Bayesian model and exact Bayesian updating, $p_{\infty}(z)$ should be independent of $\pi$ as long as $\pi$ produces dense samples across the input space (Kleijn & van der Vaart,, 2012). However, the requirements for this are very strict, with any model misspecification or error in belief updating reintroducing the dependency.